LASCO's doors have reopened following momentum management and station keeping.
LASCO doors closed until ~19:30 UT Normal C2 and C3 observations resume ~21:00 UT 21:20 UT Partial Wide bright arcade of loop fronts first observed Halo ~15Rs into C3, spanning ~150 deg. EIT CME on/over W Limb
07:27 UT S West Fairly fast loop front 12:50 UT E Limb System of very faint loop fronts, quite wide, one mainly spanning NE and other mainly SE 19:50 UT S West Narrow ragged front
01:27 UT N West Narrow ragged front along streamer 21:16 UT W Limb Very slow system of loops, quite faint and narrow
03:50 UT N East Slow wide loop front, quite bright, further similar fronts follow 09:26 UT S East Bright loop front with cavity and core 09:50 UT S East Similar to previous event but faster, brighter, narrower and further S 10:26 UT N West Bright asymmetric front along streamer, some continued outflow 20:06 UT S West Bright loop front with cavity and further loops/loop fronts 22:06 UT S East Wide bright blob with cavity
06:26 UT N West Narrow bright front with fainter extension over W Limb 08:20 - 13:45 UT Temporary data gap